OpenStudy (anonymous):

if x-y=4 and x-2y=6, what is x+y? Please help ! It's easy , I know , but can't get the answer(

OpenStudy (the_fizicx99):

Lol i was mid way through when I see it says "x+y" ._. Wonder if this helps, y = x - 4 x -2(x-4) = 6 x -2x + 8 = 6 -x + 8 = 6 -x = -2 /-x /-x x = 2 2 - y = 4 -y = 2 /-1 /-1 y = -2 2 -(-2) = 4 2 - 2(-2) = 6 2 + 4 = 6

OpenStudy (the_fizicx99):

Sorry lol I was doing mental math so I was ahead of what I was writing, 2 -(-2) = 6 2 - 2(-2) = 6 2 + 4 = 6

OpenStudy (the_fizicx99):

Lol, you understand what I did? I did this in notepad to test my theory. So its poorly written. What grade is this for btw?

OpenStudy (anonymous):

thanks too much :)))